OpenStudy (anonymous):

using induction, prove that for all integers n > 0, n3 + 2n is divisible by 3. please help.

OpenStudy (anonymous):

first of all maybe we can see why it is true without induction

OpenStudy (anonymous):

or just do it by induction either way have you ever done a proof by induction?

OpenStudy (anonymous):

i'm supposed to solve it using induction. it's obvious that it's true. even if i plug in 0 it works. but, i don't understand understand how to set this up. i know i need a base case. would that be n=0?

OpenStudy (anonymous):

no, base case is \(n=1\)

OpenStudy (anonymous):

in which case you get 3, so done with the base case

OpenStudy (anonymous):

next you assume it is true if \(n=k\) in other words you say "we assume \(k^3+2k\) is divisible by \(3\)" and using that, prove it is true if \(n=k+1\)

OpenStudy (anonymous):

most of it now is algebra from here on in replace \(n\) by \(k+1\) and separate out a part of the result that looks like \(k^3+2k\) and the rest will be obviously divisible by \(3\)

OpenStudy (anonymous):

one sec let me try to absorb what you said. thank you for helping me. i will reply in a min with questions

OpenStudy (anonymous):

take your time, do the algebra

OpenStudy (anonymous):

what is the significance of substituting n for k?

OpenStudy (anonymous):

you substitute \(k\) for \(n\) to say what you get to assume is true

OpenStudy (anonymous):

not \(n\) for \(k\)

OpenStudy (anonymous):

ok, so by saying k we are just saying we assume it's true. this signifies induction?

OpenStudy (anonymous):

and if k+1 is true then the proof is complete?

OpenStudy (anonymous):

yes that is the "induction hypothesis" assume it is true if \(n=k\) then you use that to prove it is true if \(n=k+1\)

OpenStudy (anonymous):

some people cheat and assume it is true for \(n\) then show it is true for \(n+1\) but really we should change the variable

OpenStudy (anonymous):

are you working on the algebra?

OpenStudy (anonymous):

i am trying to work on the algebra. this is what i have to solve: (k+1)^3+2(k+1)=3

OpenStudy (anonymous):

oh no not \(=3\) no one says it is equal to 3

OpenStudy (anonymous):

just take \[(k+1)^2+2(k+1)\] and expand it, then take out the part that looks like \(k^3+2k\) and separate it from the rest

OpenStudy (anonymous):

oops i meant \[(k+1)^3+2(k+1)\]

OpenStudy (anonymous):

go ahead and expand it, let me know what you get

OpenStudy (anonymous):

still doing the algebra?

OpenStudy (anonymous):

i am.. i feel retarded. thanks again for helping :\ i am trying to figure out what else i do after (k+1)((k+1)^2)+2(k+1)

OpenStudy (anonymous):

multiply it all out there is no way around it

OpenStudy (anonymous):

do you know how to cube \(k+1\) easily?

OpenStudy (anonymous):

in general \[(a+b)^3=a^2+3a^2b+3ab^2+b^2\] in your case \(a=k,, b=1\)

OpenStudy (anonymous):

and of course \(2(k+1)=2k+2\)

OpenStudy (anonymous):

okay. i have to do (k+1)(k+1)(k+1) and multiply all of it.

OpenStudy (anonymous):

yes, but i showed you how to do it in one step

OpenStudy (anonymous):

yeah i got it. sorry about that

OpenStudy (anonymous):

take your time, let me know what you get when you do all this algebra then we will be done in two steps after

OpenStudy (anonymous):

what did you mean by separate it from the rest? it's all multiplied out. k^2+3k^2(1)+3k(1)^2+1^2 + 2k+1 but there is no right side

OpenStudy (anonymous):

ok you should have \[k^3+3k^2+3k+1+2k+2\]

OpenStudy (anonymous):

you forgot to distribute with \(2(k+1)\) i think

OpenStudy (anonymous):

and also you wrote \(1^2\) but i am sure you didn't mean to

OpenStudy (anonymous):

now only two steps separate out the \(k^3+2k\) part, that we get to assume is divisible by 3

OpenStudy (anonymous):

you have \[\color{blue}{k^3+2k}+\color{red}{3k^2+3k+3}\]

OpenStudy (anonymous):

the blue part is divisible by 3 "by induction" we got to assume that the red part is divisible by 3 because... well because it is you can factor a 3 out of it if you want to be formal

OpenStudy (anonymous):

if the red wasn't divisible would it not be a proof?

OpenStudy (anonymous):

yeah you need that part to be divisible by 3 the blue part is the induction hypotheses the red part is what you have to show is divisible by 3, but there is not much to show since each term has a 3 in it

OpenStudy (anonymous):

\[\color{blue}{k^3+2k}+\color{red}{3k^2+3k+3}\] \[\overbrace{\color{blue}{k^3+2k}}^{\text{ div by 3 by induction}}+\overbrace{\color{red}{3(k^2+k+1)}}^{\text{ has a factor of 3}}\]

OpenStudy (anonymous):

and now we are done hope this helped clarify how to do a proof by induction they all pretty much work this way

OpenStudy (anonymous):

i can't believe i found someone to walk me through that

OpenStudy (anonymous):

yeah it is usually a thankless task because most people have no idea about what it means so you have to explain every step of the way, and they get frustrated really it is something you have to see done a few times to understand the logic glad it helped

OpenStudy (anonymous):

if i can find out how to donate to you i will :)